Two loudspeakers are 25.0 cm apart and produce sound waves in phase. A point on the second nodal line is1.12 m from one source and 2.12 m from the other. What is the wavelength of the waves?


1
Expert's answer
2022-05-16T13:40:58-0400
"\Delta=l_2-l_1=2.12\:\rm m-1.12 \: m=1.0\:\rm m"

"\Delta=k\lambda"

"\lambda=\frac{\Delta}{k}=\frac{1.0\:\rm m}{2}=0.5\:\rm m=50\: cm"